

********************************************************************
UARS WEEKLY STATUS REPORT - 25 MAY 2001
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, taking solar data on all available
orbital tracking periods and stellar calibration observations whenever
possible. We continue to gather enough data to meet all the SOLSTICE
science objectives during this period of reduced tape recorder operations.
***********
SUSIM
***********
SUSIM performed its daily solar scans for this past week during
periods of predicted UARS telemetry.
The file containing the SUSIM V19r3 MgII index, available from the
SUSIM WWW site (http://wwwsolar.nrl.navy.mil/) has been updated to
include an additional field, namely, the fraction of the day after
UTC midnight identifying the time of the MgII index measurement.
This data has become more significant recently during the present
solar maximum, a time period when the solar UV irradiance changes
occur more quickly.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
HRDI continues to function normally, making wind measurements during
the sunlit parts of orbits nearly every day, and measuring nighttime winds
when power is available.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Daily processing of instrument Daily Activity Plans
continues with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 12 May 2001 (Orbit 52740, GMT Day 132) through 18 May 2001
(Orbit 52845, GMT Day 138).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 135/2347 52905 Power ON
PEM AXIS 136/1222 52913 Power ON
The beta angle ranged between the angles of 70.6 degrees to 48.8 degrees for
this report period. The beta angle is now decreasing towards a minimum angle
of 34 degrees on 18 June 2001 (DOY 165).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 1.0 to 1.3 degrees C. Additional battery performance data for each
battery on 11 May 2001 and 18 May 2001 are provided in Appendix D.
The clock error ranged from Ð12.2 to 11.3 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
¥ OUT-OF-LIMITS = 00
¥ ALERTS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
¥ FOT = 14
¥ PEM = 02
¥ HALOE = 01
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
¥ FOT = 06
Six (6) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
· Transponder A: 00/316 - 2287.496493
· Transponder B: 01/119 - 2287.499116
GROUND SYSTEM OPERATION
Three (6) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 01/137 show the observatory data loss to be 300 hours, 38
minutes, 10.288 seconds. This shows an increase of 14 minutes and 58.072
seconds since the last report period. This is a 0.39 percent data loss, which
equals a 99.61 percent data capture for the mission.
The increase of 0.864 seconds occurred over one (1) day:
· 1 minute and 31.936 seconds on Day 132 due to a Late AOS TTR#14170.
· 37.632 seconds on Day 133 due to a Data Hit
· 8 minutes and 43.808 seconds on Day 134 due to weather degradation at
White Sands
· 1 minute and 7.696 seconds on Day 135 due to a Data Hit
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The new flight software patch uplinked on 25 April 2001 continues to be
monitored this week. The FOT is pleased with its overall performance and is
fine-tuning the TAM alignment by loading new alignment matrices. This is a
slow process and it will take many alignments to lock in on the most accurate
solution.
An instrument anomaly was experience on 05 May 2001. The instrument involved
was the ACRIM II instrument. All ACRIM telemetry for currents and voltages all
dropped to zero values suddenly. The ACRIM team in coordination with the
UARSFOT is conducting analysis and command tests to determine the cause of the
anomaly. The reason for the anomaly is currently unknown.
UARS is now through its high beta period in which the FSS no longer has view of
the sun. The attitude remained solid and the first sun update showed residuals
of approximately 300 arc seconds on 14 May 2001.
MEETINGS
The FOT attended a FORMATS Meeting on Tuesday, May 15, 2001.
The FOT attended an OIRB Meeting on Tuesday, May 15, 2001.
The FOT attended a UARS TAC RS Meeting on Wednesday, May 16, 2001.
The FOT attended an ERBS ORR Meeting on Thursday, May 17, 2001.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 30 May 2001 (DOY 150). This maneuver
will be a Forward to Reverse Flight maneuver (Northern Hemisphere viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 18 May 2001
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
52870 133/1507 01-302 OBC Clk Adj 20.6 to 41.2 msec/day
52887 134/1910 01-303 OBC Clk Adj 41.2 to 20.6 msec/day
52894 135/0629 01-304 OBC Clk Adj 20.6 to 0.0 msec/day
52901 135/1731 01-306 OBC TABLE 65 (STAR CATALOG)
52902 135/1830 01-307 NBTR NOOPS D-136
52904 135/2203 01-305 PEM PEM TO DAY-ONLY OPS
52905 135/2347 01-308 HALOE Power On
52911 136/0855 01-307 NBTR NOOPS D-136
52913 136/1222 01-309 PEM AXIS Power ON
52917 136/1831 01-310 NBTR NOOPS D-137
52924 137/0632 01-311 OBC Clk Adj 0.0 to 20.6 msec/day
52926 137/0857 01-310 NBTR NOOPS D-137
52932 137/1832 01-312 NBTR NOOPS D-138
52940 138/0712 01-312 NBTR NOOPS D138
52944 138/1411 01-313 ACS TAM Alignment
52947 138/1834 01-315 NBTR NOOPS D139
52949 138/2208 01-314 OBC Clk Adj 20.6 to 0.00 msec/day
____________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 18 May 2001
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period
_______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 18 May 2001
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
52870 133/1502 Clk Adj 20.6 to 41.2 msec/day
52887 134/1912 Clk Adj 41.2 to 20.6 msec/day
52894 135/0632 Clk Adj 20.6 to 0.0 msec/day
52924 137/0637 Clk Adj 0.0 to 20.6 msec/day
52949 138/2212 Clk Adj 20.6 to 0.00 msec/day
_______________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 18 May 2001
11 May 2001 - GMT Day 131 Beta = 72.9 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- -------- ---------- ------ ------- ------
1 +728.0/+728.0 -2.42/-3.99 +0.0/+0.0 11.8 0.000 0.0
2 +00.0/-28.0 +3.23/+1.97 +16.8/-16.0 26.7 1.020 15.6
3 +16.8/-22.4 +2.60/+1.34 +16.0/-16.0 26.7 1.010 15.3
18 May 2001 - GMT Day 138 Beta = 48.8 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- -------- ---------- ------ ------- ------
1 +728.0/+728.0 -3.68/-4.94 +0.0/+0.0 12.0 0.000 0.0
2 +05.6/-78.4 +4.51/+3.23 +16.8/-16.0 25.4 1.030 20.5
3 +22.4/-22.4 +3.55/+2.60 +16.0/-16.0 25.4 1.020 20.5
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
---- ---- ----------
-no changes V/T 5
_______________________________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 18 May 2001
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-025 FOT 52860 132/2343 14170 Generic Late Acquisition
01-026 FOT 52877 134/0205 23876 Late Acq. due to weather at
STGT
01-025 FOT 52917 136/1831 14170 Generic Late Acquisition
01-027 FOT 52928 137/1225 23880 Degraded Data
01-028 FOT 52931 137/1732 23880 Degraded Data
01-029 FOT 52947 138/1913 23882 Degraded 32K Data
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-026 FOT 52877 134/0205 23876 Late Acq. due to weather at
STGT
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-025 FOT see list May 01 14170 Generic Late Acquisition
01-022 ACRIM 52709 122/2140 N/A Red Low voltage alert
01-024 ACRIM 52753 125/2045 N/A ACRIM Alerts - Power Off
01-027 FOT 52928 137/1225 23880 Degraded Data
01-028 FOT 52931 137/1732 23880 Degraded Data
01-029 FOT 52947 138/1913 23882 Degraded 32K Data
SEVNTFW01/20